OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 9 white balls and 9 yellow balls and the second urn contains 8 white balls and 3 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

1/2*8/11

OpenStudy (anonymous):

huh ??

OpenStudy (anonymous):

Okay so you need to figure out the probability that both of the balls are white. In urn A, the probability that the ball you pick out will be white is 9/18. We can show this in probability as P(A) = 9/18 = 1/2 In urn B, the probability that the ball you pick out will be white is 8/11. We can show this in probability as P(B) = 8/11. Then, to find the probability of both events happening at the same time, you multiply the probabilities together which is why I had 1/2*8/11. The answer is 8/22 = 36.363636%